OpenStudy (anonymous):

, Find f'(x) if f(x)=sin^3(4x)

OpenStudy (anonymous):

f(x)=1/4*{4sin^3(4x)} =1/4*(3sin4x-sin12x) f'(x) =1/4*(12cos 4x-12cos 12x

OpenStudy (lalaly):

no its f'(x)= 3cos(4x) *4 = 12cos(4x)